2 children are balanced on a seesaw. One of them weighing 40 lbs is 6 feet from the fulcrum. How far from the fulcrum on the other side must the other child sit if he weighs 60 lbs?

Answer by stanbon(75887)

2 children are balanced on a seesaw. One of them weighing 40 lbs is 6 feet from the fulcrum. How far from the fulcrum on the other side must the other child sit if he weighs 60 lbs?
------------------------
Equation:
40*6 = 60x
-----
x = (40*6)/60
---
x = 40/10
--
x = 4 ft.
===================
Cheers,
Stan H.